OpenStudy (anonymous):

find the derivative: y =(2x-7)^3 and y= (3x^2+1)^4

OpenStudy (anonymous):

use power rule along with chain rule

OpenStudy (anonymous):

hmm so for the first example, we do this?

OpenStudy (calculusfunctions):

\[y =(2x -7)^{3}\] \[\frac{ dy }{ dx }=3(2x -7)^{2}(2)\] \[\frac{ dy }{ dx }=6(2x -7)^{2}\]

OpenStudy (anonymous):

what was that formula? i got up to 3(2x-7)^2

OpenStudy (anonymous):

Use chain rule, you still need to derive what is in the parentheses.

OpenStudy (calculusfunctions):

\[y =(3x ^{2}+1)^{4}\] \[\frac{ dy }{ dx }=4(3x ^{2}+1)^{3}(6x)\] \[\frac{ dy }{ dx }=24x(3x ^{2}+1)^{3}\]

OpenStudy (calculusfunctions):

Did you understand?

OpenStudy (anonymous):

dy/dx=3(2x−7)^2 (2)

OpenStudy (anonymous):

i look at that and i understand where the 3 comes from, but not the 2 on the right

OpenStudy (calculusfunctions):

The rule you're using is called the power of a function rule (a form of the chain rule). If\[y =[f(x)]^{n}\]then\[\frac{ dy }{ dx }=n[f(x)]^{(n -1)}[f \prime(x)]\]In other words, multiply the power (exponent) of the function by the number in the front, reduce the power of the function by one, and multiply by the derivative of the function. In other words, multiply the exponent that the bracket is raised to by the m=number in front of the bracket, then reduce the exponent of the bracket by one, then multiply by the derivative of the function inside the bracket. Kapish?!

OpenStudy (calculusfunctions):

m=number is just a typo so ignore that. it's just 'number'. OK?

OpenStudy (calculusfunctions):

So do you understand? Comprendre?

OpenStudy (anonymous):

so for example, 3(9x-4)^4 would be 4(3)(9x-4)^3 (9)?

OpenStudy (calculusfunctions):

Excellent!

OpenStudy (anonymous):

yay :D

OpenStudy (anonymous):

so i take the 4, multiply it with the coefficient outside of the parenthesis, subtract 1 from the exponent, find the derivative of 9x-4 which is just 9. gotcha.

OpenStudy (calculusfunctions):

Beautiful! Just simplify by multiplying all the monomials so for example your simplified answer should be\[\frac{ dy }{ dx }=108(9x -4)^{3}\] in your example. Alright?

OpenStudy (anonymous):

so for 2(x^2-1)^3 i would get 3*2(x^2-1)^2(1)

OpenStudy (anonymous):

yep, sure did. i was just trying to nail the basic part of it

OpenStudy (anonymous):

Well you could do this without the chain rule if you haven't learnt it but it would take a while to expand that expression.

OpenStudy (calculusfunctions):

Perfect!

OpenStudy (anonymous):

hmmmm, i have this next problem y=1/(x-2) here's what i did

OpenStudy (anonymous):

For that you can use the chain rule or the quotient rule.

OpenStudy (calculusfunctions):

Rewrite as \[y =(x -2)^{-1}\]and perform the same magic.

OpenStudy (anonymous):

yea, i started off doing that

OpenStudy (anonymous):

y=-(1/(x-2)^2)

OpenStudy (calculusfunctions):

Simply express your final answer using positive exponents. Do you want to show me your solution so I can check?

OpenStudy (anonymous):

Or you could use the quotient rule: \[\frac{ (x-2)\frac{ d }{ dx }(1)-(1)\frac{ d }{ dx } (x-2)}{(x-2)^2 }\]

OpenStudy (calculusfunctions):

OH! Yes, that is correct!

OpenStudy (anonymous):

But what Calculus functions said is a LOT easier.

OpenStudy (anonymous):

oh my gosh, i must've done about 15 quotient problems today. i CAN'T STAND THOSE ANYMORE!!!

OpenStudy (anonymous):

Haha...

OpenStudy (calculusfunctions):

Yes you could use the quotient rule, but it's not necessary.

OpenStudy (anonymous):

i got really hard ones involving square roots and such. ruined my friday :(

OpenStudy (anonymous):

Chain rule is your friend for square roots.

OpenStudy (anonymous):

i'll post a question about the ones i was having trouble with. i still didn't finish 1 problem lol

OpenStudy (anonymous):

Go ahead.

OpenStudy (calculusfunctions):

Although remember one thing. The only reason quotient rule is not necessary is because the numerator is a constant.

OpenStudy (calculusfunctions):

Are you posting them now? Because I have to log off in a few minutes. If I don't see them and you want my help then I'll help the next time I log on. Just message me to remind me.

OpenStudy (anonymous):

I can help out for a while if you were to log out...

OpenStudy (calculusfunctions):

Do you want to give me one or two quickly right now before I log out?

OpenStudy (anonymous):

sure, let me type it out really quick.

OpenStudy (anonymous):

differentiate the function: \[f(x)=\sqrt[3]{x}(\sqrt{x}+3)\]

OpenStudy (anonymous):

You have to combine the chain rule and the product rule. Or you can expand it out.

OpenStudy (anonymous):

so basically what i've learned is to find the primes and then set it up like f*g' - f' *g / (g)^2

OpenStudy (anonymous):

unless i'm using the wrong formula lol

OpenStudy (anonymous):

Expanding it will be simpler I believe.

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

So that's the same thing as \[x^\frac{ 1 }{ 3 }(x^\frac{ 1 }{ 2 }+3)\]

OpenStudy (anonymous):

i wrote it down like this: f prime = 1/3x^-1/6

OpenStudy (anonymous):

Which is the same as: \[x^\frac{ 5 }{ 6 }+3x^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

oops i meant g prime is x^1/2

OpenStudy (anonymous):

Ohh you are using the product rule. It's a LOT easier to expand and use the sum rule instead.

OpenStudy (anonymous):

yea, the way my teacher is doing it is just breaking it down into mini steps so we get a basic idea of the rules and such. to see how it all fits in. i do get what you mean though

OpenStudy (anonymous):

So the derivative is just: \[\frac{ 5 }{ 6 }x^\frac{ -1 }{ 6 } +\frac{ 3 }{ 2 }x^\frac{ -1 }{ 2}\]

OpenStudy (anonymous):

I used the power rule and the sum rule.

OpenStudy (calculusfunctions):

I'm not sure who's answering here. Do you want my explanation or are you good?

OpenStudy (anonymous):

Ohh. Sorry @calculusfunctions .

OpenStudy (calculusfunctions):

No @Dido525, absolutely no need to apologize. I just wasn't sure if she needs me stick around to help or should I log out.

OpenStudy (anonymous):

so that's all we had to do dido? lol. seems like a no brainer

OpenStudy (anonymous):

Yeah it's pretty easy.

OpenStudy (anonymous):

You can use the product rule, but that was a LOT faster.

OpenStudy (anonymous):

i was writing it originally as :

OpenStudy (anonymous):

\[(1/3x ^{-1/6})(\sqrt{x}+3)-(\sqrt[3]{x})(x ^{-1/2}) /(\sqrt{x} +3)^2\]

OpenStudy (calculusfunctions):

\[y =\sqrt[3]{x}(\sqrt{x}+3)\] \[y=x ^{\frac{ 1 }{ 3 }}(x ^{\frac{ 1 }{ 2 }}+3)\] \[y =x ^{\frac{ 5 }{ 6 }}+3x ^{\frac{ 1 }{ 3 }}\] \[\frac{ dy }{ dx }=\frac{ 5 }{ 6 }x ^{\frac{ -1 }{ 6 }}+\frac{ 1 }{ 3 }x ^{\frac{ -2 }{ 3 }}\] \[\frac{ dy }{ dx }=\frac{ 5 }{ 6\sqrt[6]{x} }+\frac{ 1 }{ 3\sqrt[3]{x ^{2}} }\] You can also use the product rule but it's not necessary when the exponent raised to the bracket is just one.

OpenStudy (anonymous):

i put a - in x^-1/2 by accident lol. im so tired

OpenStudy (calculusfunctions):

@Dido525, I am truly sorry as it was not my intention to offend you or be rude. I sincerely apologize if I came across as snobbish or insensitive. That's not who I am.

OpenStudy (anonymous):

@calculusfunctions . You are crazy!!! I never once thought that!! AHAH!!! Lol, you made my day :P .

OpenStudy (anonymous):

o.0

OpenStudy (calculusfunctions):

@Dido525, good! I'm glad to hear that because I felt terrible. Thanks.

OpenStudy (anonymous):

now we can all go to sleep happy :D

OpenStudy (calculusfunctions):

@qtexpress101, do you understand now? I'm sure @Dido will do an excellent job at explaining when I log out.

OpenStudy (anonymous):

yea, i have a better understanding thanks to the both of you. really appreciate it!

OpenStudy (calculusfunctions):

Welcome and Goodnight!

OpenStudy (anonymous):

goodnight!